# Learning from Imperfect Human Feedback: a Tale from Corruption-Robust $\text{Dueling}^{*}$

### v1

@@ B.2 Proof for Proposition 1: Regret Lower Bound for Linear Utility
There exists an LIHF instance with a $\rho$-imperfect user and linear user utility such that any learner knowing $\rho$ has to suffer $\text{Reg}_{T}\geq\Omega(d\max\{\sqrt{T},T^{\rho}\})$.
Proof  (Proof for Proposition 1). The proof structure is very similar to the lower bound proof in Theorem 1. We start by discussing the value of $\rho$. Consider the scenario when $\rho\leq 0.5$. Applying Lemma B.6 and B.1 together with linear link function, we have $\text{Reg}_{T}\geq\Omega(d\max\{\sqrt{T},T^{\rho}\})$, which completes the proof.
Lemma B.6 ((Lattimore and Szepesvári, 2020)).
Let $\mathcal{A}=[-1,1]^{d}$ and $\Theta=[-T^{-\frac{1}{2}},T^{-\frac{1}{2}}]^{d}$. Consider the linear reward function $r_{t}=\theta^{\top}A_{t}+\epsilon_{t}$, $\epsilon_{t}$ is independent Gaussian noise with mean 0 variance 1. Then for any learner $\mathcal{L}$, there exists a vector $\theta\in\Theta$ such that
@@
Now, let’s focus on the scenario when $\rho>0.5$ and the essence is to extend Lemma B.6 to the scenario in presence of adversarial corruption induced by $\rho$-imperfect user, which is shown in Lemma B.7
Lemma B.7.
Assume that the action space $\mathcal{A}\subset[-1,1]^{d}$. Given corruption induced by $\rho$-imperfect user (Def. 1), for stochastic linear bandit, there exists a $\theta\in[-C_{\kappa}T^{\rho-1},C_{\kappa}T^{\rho-1}]^{d}$ such that any learner $\mathcal{L}$ even with the knowledge of $\rho$ has to suffer regret no less than $\frac{d}{8}C_{\kappa}T^{\rho}.$
Proof  (Proof for Lemma B.7). The proof extends Lemma B.6 to a scenario in presence of corruption. We want to construct a parameter family and a corruption strategy such that for all algorithm $\mathcal{L}$, it will occur at least $\Omega(dT^{\rho})$ regret. Consider the action set $\mathcal{A}\in[-1,1]^{d}$ and $\Theta=\{-\beta,\beta\}^{d}$. For any learner $\mathcal{L}$, we can lower bound its regret by
@@
Let $i\in[d]$ and $\theta\in\Theta$ be fixed, and let $\theta^{\prime}_{j}=\theta_{j}$, for $j\neq i$ and $\theta^{\prime}_{i}=-\theta_{i}$. Then using Lemma B.8, we have
$p_{\theta_{i}}+p_{\theta^{\prime}_{i}}\geq\frac{1}{2}\exp(-\mathbb{D}(\mathbb{P}_{\theta},\mathbb{P}_{\theta^{\prime}}))=\frac{1}{2}\exp\left(-\mathbb{E}\left[\sum^{T}_{t=1}\text{D}_{\text{KL}}(P_{a_{t}},P^{\prime}_{a_{t}})\right]\right).$
$P_{a_{t}}$ is the distribution of corrupted reward observed by $\mathcal{L}$ after playing action $a_{t}$ when reward parameter is $\theta$. Similarly, $P^{\prime}_{a_{t}}$ is the distribution of corrupted observed by $\mathcal{L}$ after playing action $a_{t}$ when the reward parameter is $\theta^{\prime}$.
Lemma B.8 (Bretagnolle-Huber inequality).
@@
Consider the corruption strategy such that $c_{t}(a_{t}|\theta)=-a_{t,i}\theta_{i}$ in (11) and $c_{t}(a_{t}|\theta^{\prime})=a_{t,i}\theta_{i}$ in (12). This is achievable since we assume the adversary has complete knowledge of the problem instance. By doing so, the corrupted reward $\hat{r}_{t}$ are from the same distribution regardless whether the preference parameter is $\theta$ or $\theta^{\prime}$. This is because $\theta$ and $\theta^{\prime}$ only differs in the $i$-th coordinate with the magnitude $\beta$, and this difference could be masked by the corruption, resulting in $\mathbb{E}\left[\sum^{T}_{t=1}\mathbb{D}(P_{A_{t}},P^{\prime}_{A_{t}})\right]=0$, which makes it indistinguishable by $\mathcal{L}$. Notice that executing such a corruption strategy requires total corruption budget
@@
which completes the proof. We want to highlight that the corruption level each round $|c_{t}(a_{t})|$ is bounded by $\beta\|a_{t}\|_{\infty}\leq C_{\kappa}T^{\rho-1}\leq C_{\kappa}t^{\rho-1},\forall t$, satisfying the corruption induced by $\rho$-imperfect user.
C Missing Proofs in Theorem 2 and Proposition 2
C.1 Proof for Theorem 2: Efficiency-Robustness Tradeoff in DBGD
In this section, we present the regret upper for dueling bandit gradient descent Yue and Joachims (2009) in presence of arbitrary and agnostic corruption. Before presenting the proof, we make several remarks.
First, in Yue and Joachims (2009), the utility function $\mu$ is assumed to be strictly concave to ensure the uniqueness of the maximizer. However, in scenarios where the maximizer is unique within the action set for a concave utility function $\mu$, the requirement for ”strictness” can be relaxed.
Second, we highlight the notation differences. We use $a$ to denote actions while Yue and Joachims (2009) uses $w$. In our proof, we define $P_{t}(a):=\sigma\left(\mu(a_{t})-\mu(a)\right)$ to represent the probability of the event $a_{t}\succ a$. While in Yue and Joachims (2009), it is denoted as $c_{t}(w)$. Just to highlight, its corrupted version $\hat{P}_{t}(a):=\sigma\left(\mu(a_{t})-\mu(a)+c_{t}(a_{t},a)\right)$ is newly introduced in our paper. Moreover, $\bar{P}_{t}(a):=\mathbb{E}_{x\in\mathbb{B}}\left[P_{t}(\mathcal{P}_{\mathcal{A}}(a+\delta x))\right]$ is denoted as $\hat{\epsilon}_{t}(w)+\frac{1}{2}$ in Yue and Joachims (2009). The definition of $\mathcal{F}\left(\mathcal{P}_{\mathcal{A}}\left(a_{t}+\delta u_{t}\right),a_{t}\right)$ is same as $X_{t}\left(\mathcal{P}_{W}(w_{t}+\delta u_{t})\right)$ defined in Yue and Joachims (2009), except different notation.
@@ C.1 Proof for Theorem 2: Efficiency-Robustness Tradeoff in DBGD
For any $\alpha\in(0,\frac{1}{4}]$ and any number of round $T$ satisfying $T>\left(\frac{\sqrt{2Rd}L_{\mu}L_{2}}{\sqrt{13L}L_{\sigma}}\right)^{4}$, Algorithm 1 parameterized with $\gamma=\frac{R}{\sqrt{T}}$ and $\delta=\frac{\sqrt{2Rd}}{\sqrt{13L}T^{\alpha}}$ achieves $\text{Reg}_{T}\leq O(\sqrt{d}T^{1-\alpha}+\sqrt{d}T^{\alpha}C)$ for any total corruption level $C$.
Proof  (Proof for Theorem 2). The proof of Theorem 2 builds on the proof of Yue and Joachims (2009), where we extend it to a setting in presence of agnostic corruption. The cornerstone of the theoretical analysis relies on developing the regret decomposition lemma (Lemma 2), where we novelly decompose the regret into two parts: regret of decision and observation error. This decomposition is achieved through quantifying the bias, $b_{t}$, caused by adversarial corruption, in gradient estimation (Lemma C.1). We further use Lemma C.5 to upper bound the observation error in terms of the exploration size $\delta$. Lemma 2 and Lemma C.5 reveal that the regret of decision and observation error are jointly controlled by the exploration size $\delta$. We choose $\delta$ in order of $O(T^{-\alpha})$ to achieve a balance between optimizing the learning objective (1) and controlling the gradient bias arising from the corrupted dueling feedback. Consequently, we have $\text{Reg}_{T}$ is upper bounded by $O(\sqrt{d}T^{1-\alpha}+\sqrt{d}T^{\alpha}C)$.
The development of the proof is based on the observation that we decompose bias in the gradient caused by corruption. Notice that without corruption, DBGD performs expected gradient ascent, where the gradient $g_{t}$ is defined as follows
In the presence of adversarial corruption, the corrupted gradient is
Therefore, we quantify the bias in the following lemma.
C.1.1 Proof for Lemma C.1: Gradient Estimation Under Corruption
Lemma C.1.
Proof  (Proof for Lemma C.1). Since $a^{\prime}_{t}:=\mathcal{P}_{\mathcal{A}}(a_{t}+\delta u_{t})$, we have
We get the equation (14) by using Lemma C.2. We get equation (15) by using Lemma C.3.
Lemma C.2 (Lemma 2 in Yue and Joachims (2009)).
Fix $\delta>0$, over random unit vector $u$, we have
Lemma C.3 (Lemma 1 in Yue and Joachims (2009)).
C.1.2 Proof for Lemma 2: Regret Decomposition
Lemma (Lemma 2 restated).
Choose $\lambda=\frac{L_{\sigma}}{L_{\sigma}-\delta L_{\mu}L_{\sigma}},\gamma=R/\sqrt{T}$, for $b_{t}$ defined in Lemma C.1, we have
[…]
Proof  (Proof for Lemma 2). Because of Lemma C.4, we have
$\mathbb{E}\left[\sum^{T}_{t=1}\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*})\right]\leq\sum^{T}_{t=1}\mathbb{E}\left(\lambda\nabla\bar{P}_{t}(a_{t})(a_{t}-a^{*})+(3+\lambda)\delta L\right)$
$=\lambda\sum^{T}_{t=1}\mathbb{E}(\mathbb{E}(g_{t}|a_{t})^{\top}(a_{t}-a^{*}))+(3+\lambda)\delta LT$
$=\lambda\sum^{T}_{t=1}\mathbb{E}(\mathbb{E}(\hat{g}_{t}+b_{t}|a_{t})^{\top}(a_{t}-a^{*}))+(3+\lambda)\delta LT$
$=\lambda\sum^{T}_{t=1}\mathbb{E}(\hat{g}_{t}^{\top}(a_{t}-a^{*}))+\lambda\sum^{T}_{t=1}\mathbb{E}(b_{t}^{\top}(a_{t}-a^{*}))+(3+\lambda)\delta LT$
Lemma C.4 (Lemma 4 in Yue and Joachims (2009)).
Fix $\delta\in(0,\frac{L_{\sigma}}{L_{\nu}L_{\sigma}})$ and define $\lambda=\frac{L_{\sigma}}{L_{\sigma}-\delta L_{\nu}L_{2}}$, we have
$\mathbb{E}\left[\sum^{T}_{t=1}\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*})\right]\leq\sum^{T}_{t=1}\mathbb{E}\left(\lambda\nabla\bar{P}_{t}(a_{t})^{\top}(a_{t}-a^{*})+(3+\lambda)\delta L\right).$
Noticing that $\|b_{t}\|_{2}\leq\frac{d}{\delta}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)$. Moreover, $\|a_{t}-a^{*}\|_{2}\leq 2R$. Therefore, we can control the observation error by the following.
C.1.3 Proof for Lemma C.5: Observation Error
Lemma C.5.
$\sum^{T}_{t=1}\mathbb{E}\left(b_{t}^{\top}(a_{t}-a^{*})\right)\leq 2RdL_{\sigma}C/\delta$.
Proof  (Proof for Lemma C.5).
$\sum^{T}_{t=1}\mathbb{E}(b_{t}(a_{t}-a^{*}))$ $\leq 2R\frac{d}{\delta}\sum^{T}_{t=1}\min\left(2,L_{\sigma}|c_{t}(a_{t},a^{\prime}_{t})|\right)$
$\leq 2R\frac{dL_{\sigma}}{\delta}\sum^{T}_{t=1}|c_{t}(a_{t},a^{\prime}_{t})|$
$=2R\frac{dL_{\sigma}C}{\delta}.$
Lemma C.6 (Lemma 3 in Yue and Joachims (2009)).
$\text{Reg}_{T}\leq 2\mathbb{E}\left[\sum^{T}_{t=1}\bar{P}_{t}(a_{t})-\bar{P}_{t}(a^{*})\right]+5\delta LT$
Set $G=\frac{d}{\delta}$, and by Lemma C.6, we have
$\text{Reg}_{T}\leq\lambda\left(\frac{2Rd\sqrt{T}}{\delta}+13\delta LT\right)+4R\frac{\lambda dL_{\sigma}C}{\delta}.$
Therefore, by setting $\delta=\frac{\sqrt{2Rd}}{\sqrt{13L}T^{\alpha}},\gamma=\frac{R}{\sqrt{T}}$, and $T>\left(\frac{\sqrt{2Rd}L_{v}L_{2}}{\sqrt{13L}L_{\sigma}}\right)^{4}$, we have
$\text{Reg}_{T}\leq 2\lambda_{T}\sqrt{26RdL}T^{1-\alpha}+2L_{\sigma}\sqrt{26RdL}T^{\alpha}C.$
$\lambda_{T}=\frac{L_{\sigma}\sqrt{13L}T^{\alpha}}{L_{\sigma}\sqrt{13L}T^{\alpha}-L_{v}L_{2}\sqrt{2Rd}}$.
C.2 Proof for Corollary 1: Minimax Lower Bound for DBGD
Theorem 2 implies that DBGD could afford agnostic corruption level $O(T^{3/4})$. This implies that there exists a hard instance which makes DBGD attain regret in order of $\Omega(T^{3/4})$, formally described as follows.
Corollary (Corollary 1 restated).
There exists a linear utility function $\mu$ and a link function $\sigma$ such that choosing $\alpha=1/4$, $\text{Reg}_{T}$ suffered by Algorithm 1 is at least $\Omega\left(T^{3/4}\right)$.
Proof  (Proof for Corollary 1). To start a proof by contradiction, we assume that Theorem 2 is false. Specifically, for all problem instance $\mu,\sigma$, there exist a $\epsilon>0$ such that $\text{Reg}_{T}\leq O(T^{3/4-\epsilon})$. We assume $\epsilon$ to be the least possible. In particular, there exists a pair of instance $\mu,\sigma$ such that $\text{Reg}_{T}=c_{0}T^{3/4-\epsilon}$, $c_{0}$ is a positive constant. In other words, it says that there exists a pair of instance $\mu,\sigma$ such that DBGD suffers regret in $\Theta(T^{3/4-\epsilon})$.
Consider the following problem instance. $\mu(a)=\theta^{\top}a$, specifically $\mu$ is linear function. Let $d=2$ with action set $\mathcal{A}_{1}:=\{(a_{1},a_{2}):a_{1}\geq 0,a_{2}\geq 0,\frac{1}{2}a_{1}+a_{2}-\frac{1}{4}\leq 0\}$ with $\theta=[\frac{1}{2},\frac{1}{2}]$. The optimal arm is $a_{1}=[\frac{1}{2},0]$, which is unique. Let the link function $\sigma(x)=\frac{1}{2}+\frac{1}{2}x$. It is easy to verify that it is rotational symmetric and Lipschitz. Denote $R_{T}$ as the regret occurred by DBGD on this problem instance. We know $R_{T}\leq O(T^{\frac{3}{4}-\epsilon})$. Consequently, it implies that $a_{2}$ is at most proposed $8c_{0}T^{\frac{3}{4}-\epsilon}$ times. This is because $R_{T}\leq c_{0}T^{\frac{3}{4}-\epsilon}$ and if the proposed action pair $(a_{t},a^{\prime}_{t})$ including $a_{2}$ at round $t$, it occurs regret at least $\frac{1}{8}$.
Now consider a different problem instance with the same $\mu,\sigma$ but different action set, which is $\mathcal{A}_{2}:=\{(a_{1},a_{2}):a_{1}\geq 0,a_{2}\geq 0,\frac{3}{2}a_{1}+a_{2}-\frac{3}{4}\leq 0\}$. The optimal arm in this action set $a_{2}=[0,\frac{3}{4}]$, which is also unique. Proposing arm $a_{1}=[\frac{1}{2},0]$ occurs at least $\frac{1}{8}$ regret. Consider an adversary that pulls the utility of $a_{2}$ from $\frac{3}{8}$ down to $\frac{1}{8}$ whenever it is pulled at a cost of $c_{t}(A_{t})=\frac{1}{4}$. It is equivalent to say that every time when the proposed arm $w$ falls in the region $\mathcal{W}=\{a_{1}\geq 0,a_{2}\geq 0,\frac{1}{2}a_{1}+a_{2}-\frac{1}{4}\geq 0,\frac{3}{2}a_{1}+a_{2}-\frac{3}{4}\leq 0\}$, we assume that the utility is sampled from $\theta^{\top}\mathcal{P}_{\mathcal{A}_{1}}(w)$ and $|c_{t}(w)|=|\theta^{\top}\mathcal{P}_{\mathcal{A}_{1}}(w)-\theta^{\top}w|\leq\frac{1}{4}$. Choosing $C=2c_{0}T^{\frac{3}{4}-\epsilon}$, the adversary can afford to corrupt $8c_{0}T^{\frac{3}{4}-\epsilon}$ times.
This means that as long as the adversary is corrupting, the utility observed in the first problem instance is exactly the same as the utility observed in the second problem instance, in which we pull $a_{2}$ as most $8c_{0}T^{\frac{3}{4}-\epsilon}$. However, $a_{2}$ is the optimal arm in the second problem instance, which implies that the regret occurred on the second problem instance is at least $T-8c_{0}T^{\frac{3}{4}-\epsilon}$, which is linear in $T$. This contradicts Theorem 2, which says that when agnostic corruption $C=2c_{0}T^{\frac{3}{4}-\epsilon}$, the regret upper bound is $O(T^{1-\epsilon})$, which is sublinear. To reconcile the conflicts, we must have Proposition 1 true to make the agnostic corruption $C\geq\Omega(T^{\frac{3}{4}})$ to make the parallel world argument valid.
C.3 Proof for Matching Regret Upper Bound in Theorem 1
In this section, we consider the setting when human has strongly concave utility function. When $\rho$ is known, we optimally adjusts the learning rate of the algorithm, Noisy Comparision-based Stochastic Mirror Descent (NC-SMD), proposed in Kumagai (2017), to achieve a balance between optimizing the learning objective (1) and controlling the gradient bias arising from the corrupted dueling feedback. Then we show that it attains the matching regret upper bound of $\tilde{O}(d\max\{\sqrt{T},T^{\rho}\})$ under corruption induced by $\rho$-imperfect user.
For the ease of delivering proof, in this section, we work with the cost function $f$, opposite of the utility function $\mu$, We restate the modelling of inaccuracies from human feedback and the learning objective using $f$ as follows.
Assumption C.1.
(Corrupted Utility-Based Dueling Bandit). For a pair of actions $(a,a^{\prime})\in\mathcal{A}\times\mathcal{A}$ and the corruption level $c(a,a^{\prime})$, there exists function $f:\mathcal{A}\rightarrow\mathbb{R}$ and $\sigma:\mathbb{R}\rightarrow[0,1]$ such that the probability that $a\succ a^{\prime}$ happens of receiving corrupted dueling feedback can be represented follows
^$\hat{\mathbb{P}}(a\succ a^{\prime})=\sigma(f(a^{\prime})-f(a)+c(a,a)).$
Additionally, we assume that the cost function $f$, the link function $\sigma$, and the action space $\mathcal{A}$ exhibit the following properties, which align precisely with those assumed in Kumagai (2017).
Assumption C.2.
The cost function $f:\mathcal{A}\rightarrow\mathbb{R}$ is twice continuously differentiable, $L$-Lipschitz, $\alpha$-strongly convex and $\beta$-smooth with respect to the Euclidean norm.
Definition C.1.
(Strong Convexity). A function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $\alpha$-strongly convex over the set $\mathcal{A}\subset\mathbb{R}^{d}$ if for all $x,y\in\mathcal{A}$ it holds that
$f(y)\geq f(x)+\nabla f(x)^{\top}(y-x)+\frac{\alpha}{2}\|y-x\|^{2}_{2}.$
Definition C.2.
(Smoothness). A function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $\beta$-smooth over the set $\mathcal{A}\subset\mathbb{R}^{d}$ if for all $x,y\in\mathcal{A}$ it holds that
$f(y)\leq f(x)+\nabla f(x)^{\top}(y-x)+\frac{\beta}{2}\|y-x\|^{2}_{2}.$
Assumption C.3.
Let $B:=\sup_{a,a^{\prime}\in\mathcal{A}}f(a^{\prime})-f(a)$, $f$ is defined in Assumption C.2. The link function $\sigma:\mathbb{R}\rightarrow[0,1]$ is three times differentiable and rotational-symmetric. Its first derivative is positive and monotonically non-increasing on $[0,B]$.
Assumption C.4.
We assume the action space $\mathcal{A}\in\mathbb{R}^{d}$ is convex, compact and has non-empty interior.
Definition C.3.

### v2

@@ A.2 Proof for Proposition 1:
There exists an LIHF instance with $\rho$-Imperfect Human Feedback (Definition 1), linear user utility $\mu$, and link function $\sigma$, such that any learner has to suffer $\text{Reg}_{T}\geq\Omega(d\max\{\sqrt{T},T^{\rho}\})$, even with the knowledge of $\rho$.
Proof  The proof structure is very similar to the lower bound proof in Theorem 1. We start by discussing the value of $\rho$. Consider the scenario when $\rho\leq 0.5$. Applying Lemma 12 and 7 together with linear link function, we have $\text{Reg}_{T}\geq\Omega(d\max\{\sqrt{T},T^{\rho}\})$, which completes the proof.
Lemma 12 ((Dani et al., 2008)).
Let $\mathcal{A}=[-1,1]^{d}$ and $\Theta=[-T^{-\frac{1}{2}},T^{-\frac{1}{2}}]^{d}$. Consider the linear reward function $r_{t}=\theta^{\top}A_{t}+\epsilon_{t}$, $\epsilon_{t}$ is independent standard Gaussian noise. Then for any learner, there exists a vector $\theta\in\Theta$ such that
@@
Now, let’s focus on the scenario when $\rho>0.5$ and the essence is to extend Lemma 12 to the scenario in presence of $\rho$-Imperfect Human Feedback, which is shown in Lemma 13
Lemma 13.
Assume that the action space $\mathcal{A}\subset[-1,1]^{d}$. Given $\rho$-Imperfect Human Feedback (Definition 1), for stochastic linear bandit, there exists a $\theta\in[-C_{\kappa}T^{\rho-1},C_{\kappa}T^{\rho-1}]^{d}$ such that any learner even with the knowledge of $\rho$ has to suffer regret no less than $\frac{d}{8}C_{\kappa}T^{\rho}.$
Proof  The proof extends Lemma 12 to a scenario in presence of corruption. We want to construct a parameter family and a corruption strategy such that for all algorithm, it will occur at least $\Omega(dT^{\rho})$ regret. Consider the action set $\mathcal{A}\in[-1,1]^{d}$ and $\Theta=\{-\beta,\beta\}^{d}$. For any learner, we can lower bound its regret by
@@
Let $i\in[d]$ and $\theta\in\Theta$ be fixed, and let $\theta^{\prime}_{j}=\theta_{j}$, for $j\neq i$ and $\theta^{\prime}_{i}=-\theta_{i}$. Then using Lemma 14, we have
$p_{\theta_{i}}+p_{\theta^{\prime}_{i}}\geq\frac{1}{2}\exp\left(-\mathbb{E}\left[\sum^{T}_{t=1}\text{D}_{\text{KL}}(P_{a_{t}},P^{\prime}_{a_{t}})\right]\right).$
$P_{a_{t}}$ is the distribution of corrupted reward observed by the learner after playing action $a_{t}$ when reward parameter is $\theta$. Similarly, $P^{\prime}_{a_{t}}$ is the distribution of corrupted observed by the learner after playing action $a_{t}$ when the reward parameter is $\theta^{\prime}$.
Lemma 14 (Bretagnolle-Huber inequality).
@@
Consider the corruption strategy such that $c_{t}(a_{t}|\theta)=-a_{t,i}\theta_{i}$ in (11) and $c_{t}(a_{t}|\theta^{\prime})=a_{t,i}\theta_{i}$ in (12). This is achievable since we assume the adversary has complete knowledge of the problem instance. By doing so, the corrupted reward $\hat{r}_{t}$ are from the same distribution regardless whether the preference parameter is $\theta$ or $\theta^{\prime}$. This is because $\theta$ and $\theta^{\prime}$ only differs in the $i$-th coordinate with the magnitude $\beta$, and this difference could be masked by the corruption, resulting in $\mathbb{E}\left[\sum^{T}_{t=1}D_{\text{KL}}(P_{a_{t}},P^{\prime}_{a_{t}})\right]=0$, which makes it indistinguishable by the learner. Notice that executing such a corruption strategy requires total corruption budget
@@
which completes the proof. We want to highlight that the corruption level each round $|c_{t}(a_{t})|$ is bounded by $\beta\|a_{t}\|_{\infty}\leq C_{\kappa}T^{\rho-1}\leq C_{\kappa}t^{\rho-1},\forall t$, satisfying definition of $\rho$-Imperfect Human Feedback.
B Proof for Theorem 2
@@
RoSMID satisfies $\text{Reg}_{T}\leq\tilde{\mathcal{O}}(d\max\{\sqrt{T},T^{\rho}\})$ for any $\rho$-LIHF problem.
Proof  At the beginning of the proof, we formally define self-concordance (see Definition 2). The input, $\nu$-self-concordant function $\mathcal{R}$, serves as a regularizor in RoSMID (see Algorithm 1).
Definition 2.